OpenStudy (anonymous):

write an equation of the line that passes through the given point and is PERPENDICULAR to the given line. (1, 3), y = 2x – 1

OpenStudy (anonymous):

when a line is perpendicular to another line the slope of any one line that you are trying to find the equation of is the negative reciprocal of another slope. so this means that if line 1 has equation, y=2x-1 then the slope of another perpendicular line would be -1/2 now you can make a new equation in the form y=mx+b including the new slope and then substitute the value of the coordinates you are given (1,3) and then solve for b rewrite the equation clearly showing the value of m and b

OpenStudy (radar):

Good hints.

OpenStudy (radar):

Knowing the definition of slope m and now knowing it value -1/2, the solution of the line equation is possible. Just work on it.

OpenStudy (radar):

\[m=\Delta y/\Delta x\]\[-(1/2)=\Delta y/\Delta x\]\[\Delta y =(y - 3)\]\[\Delta x=(1-x)\]